In triangle ABC , angle B = 76 degrees , side a=45cm and side b=53cm . Find angle A
Found 2 solutions by 10301650, Alan3354:
Answer by 10301650(32)

use SOH CAH TOA to solve that...
TAN A= a/b
=45/53
=0.84906
(transfer TAN to the other side)
= TAN -1 (0.84906)
=40.33325

Answer by Alan3354(69443)

In triangle ABC , angle B = 76 degrees , side a=45cm and side b=53cm . Find angle A
---------------
The other tutor seemed to assume a right triangle.
If it's not:
Use the Law of Sines.
sin(A)/a = sin(B)/b
sin(A)/45 = sin(76)/53
sin(A) = 45*sin(76)/53
sin(A) =~ 0.823836
A =~55.47 degs